OpenStudy (peterg2001):

will medal The first term of a sequence is 18, and the fifth term is 54. What is the common difference between each two consecutive terms? 36 27 18 9 @ichliebedich

OpenStudy (ichliebedich):

i think its 9 1. 18 2. 18+9=27 3. 27+9=36 4. 36+9=45 5. 45+9=54